OpenStudy (gschibby):

y^2 - x^2 - x^3 = 0 When x = 2/3

OpenStudy (anonymous):

When x = 2/3, you get y^2 - 4/9 - 8/27 = 0 y^2 -20/27 = 0 y^2 = 20/27 y = +/- sqrt (20/27)

OpenStudy (anonymous):

y =+/- 2*sqrt(15)/9

OpenStudy (gschibby):

According to my professor it's supposed to be (2*sqrt(3))/9 or (2*3^(1/2))/9 and -(2*sqrt(3))/9 or -(2*3^(1/2))/9 But hoooow? :P

OpenStudy (anonymous):

Simplifying what I wrote gives you what your professor said.

OpenStudy (anonymous):

They are all equivalent.

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

could you show us how that simplification works?

OpenStudy (anonymous):

2 * 3^1/2 does not = (2*sqrt(15))/9

OpenStudy (gschibby):

Ah damn, I wrote something wrong in the original post! It's supposed to be: y^2 - x^2 + x^3 = 0 It's a pluss sign before x^3

OpenStudy (gschibby):

Anyone? :)

OpenStudy (anonymous):

So, you get y^2 - 4/9 + 8/27 = 0 y^2 - 4/27 = 0 y = +/- sqrt (4/27) or y = +/- 2/(3 sqrt(3))

OpenStudy (anonymous):

Note: sqrt (4/27) = sqrt(4)/sqrt(27) Sqrt (4) = 2 Sqrt (27) = sqrt (9) timkes sqrt(3) = 3 sqrt (3)

OpenStudy (gschibby):

Ah! Now I get it! Thanks a lot!^^

OpenStudy (anonymous):

You're welcome!